• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 08:10:19 on June12,2010

    Entry number 325955

    keyword=Owl shift summary

    Starting the shift with no beam, on empty target.

    00:30 MCC is going to send tune beam, but there is a problem with the target ion chamber. There is an FSD fault, so they will call in the FSD expert.

    1:17 Target ion chamber is okay again; MCC sending tune beam.

    1:56 Start long beam recovery procedure.

    2:39 Done with beam recovery up to BeO running, go to thin carbon.

    Begin Q2 and try to look at cavity monitors.

    3:23 The ioc for the cavity monitors is down. Requested an escorted access so the Ahmed can reboot the cavity monitor ioc (this ioc cannot be rebooted remotely).

    3:42 ioc reboot successful (access is done)! Ahmed will now set up the gains for the cavity monitor.

    3:52 target to empty so that we can turn off the raster and set the cavity monitor gains.

    4:11 Ahmed is doing a box scan.

    The trigger rates are higher than we would expect, so we spent a little time doing spot++

    6:35 Box scan is done and still doing Q2 runs.

    Running on thin lead

    7:47 go to 50nA on thin lead

    8:00 start dipole scan

    Shift ending with dopole scan on thin lead.

    Q2 runs:

    Thin Carbon: Runs 27804/6886

    Thick Carbon: Runs 27805/6887

    Thin lead:

    1nA: Runs 27824/6906, 27825/6907, 6908, 27826/6909

    50nA: Runs 27827/6910

    Dipole scan:

    +1% dipole: Run 27830